OpenStudy (anonymous):

Vicky is twice as old as Jimmy. Jimmy is four years older than Davidl The sum of their ages is 28. how old is each person?

myininaya (myininaya):

V is twice as old as J. Do you know how to write this as an equation?

myininaya (myininaya):

is means equals by the way

OpenStudy (anonymous):

yeah i just dont know what thee formula is

myininaya (myininaya):

You do know to how write the equation? V is twice as old as J V = twice as old as J

myininaya (myininaya):

We need to figure out how to write the twice as old as J part

myininaya (myininaya):

What does twice as means?

OpenStudy (anonymous):

nevermind i got it

myininaya (myininaya):

in symbols of course lol ok

myininaya (myininaya):

What did you get then?

OpenStudy (anonymous):

x+x+4+2(x+4)=28

OpenStudy (anonymous):

the answers are 4, 8, and 16 i got it lol k thanks? ii guess hahah

myininaya (myininaya):

Ok so let me check that We have V=2J (because twice as means 2 times ) Jimmy is four years older than David J is 4 years older than David J=4 years older than David J=4+D So we have the sum of their ages is 28 That means Vicky's age+Jimmy's age+David's age=28 V + J + D =28 Now we should put this in terms of one variable so we can solve You have V=2J and D=J-4 (remember J=4+D) So Sub these into find J 2J+J+J-4=28 4J-4=28 4J=32 J=8 V=2(8)=16 D=J-4=8-4=4 Yep that is right! good job! :)